a/1-2a+b/1-2b+c/1-2c=1/2 then find the value of 1/1-2a+1/1-2b+1/1-2c.................................................................................

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
( – 1/2) [{1 – 1/(1-2a)} + {1 – 1/(1-2b)} + {1 – 1/(1-2c)}] = ½
 
{1 – 1/(1-2a)} + {1 – 1/(1-2a)} + {1 – 1/(1-2a)} =  1 + 3 = 4
 
Regards
Arun